### v1

# On Developers’ Self-Declaration of AI-Generated Code: An Analysis of Practices

Abstract.
AI code generation tools have gained significant popularity among developers, who use them to assist in software development due to their capability to generate code. Existing studies mainly explored the quality, e.g., correctness and security, of AI-generated code, while in real-world software development, the prerequisite is to distinguish AI-generated code from human-written code, which emphasizes the need to explicitly declare AI-generated code by developers. To this end, this study intends to understand the ways developers use to self-declare AI-generated code and explore the reasons why developers choose to self-declare or not. We conducted a mixed-methods study consisting of two phases. In the first phase, we mined GitHub repositories and collected 613 instances of AI-generated code snippets. In the second phase, we conducted a follow-up industrial survey, which received 111 valid responses. Our research revealed the practices followed by developers to self-declare AI-generated code. Most practitioners (76.6%) always or sometimes self-declare AI-generated code. In contrast, other practitioners (23.4%) noted that they never self-declare AI-generated code. The reasons for self-declaring AI-generated code include the need to track and monitor the code for future review and debugging, and ethical considerations. The reasons for not self-declaring AI-generated code include extensive modifications to AI-generated code and the developers’ perception that self-declaration is an unnecessary activity. We finally provided guidelines for practitioners to self-declare AI-generated code, addressing ethical and code quality concerns.
AI-Generated Code, GitHub Copilot, ChatGPT, Self-Declaration
journal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software development techniques
1. Introduction
Artificial Intelligence (AI) techniques, especially those employing Machine Learning (ML) models and Large Language Models (LLMs), are revolutionizing software engineering practices. There has been increased adoption of these models to develop more accurate, faster, and scalable tools that assist software engineers in their development tasks (Yellin, 2023) (e.g., AI-assisted coding (Liang et al., 2024)). One area that has been largely transformed with the advancement in LLMs is automatic code generation (Jiang et al., 2024; Chen et al., 2021). Developers increasingly rely on AI code generation tools in their written code (AIt, 2024; AIR, 2023). The most recent Stack Overflow Developer Survey (2024) shows that there has been a significant adoption of code generation tools in practice (SO-, [n. d.]). This is partially due to the widely available coding-specific tools (e.g., GitHub Copilot (Cop, 2024) and CodeWhisperer (Cod, 2024)) and general-purpose LLMs (e.g., ChatGPT (Cha, 2024) and Gemini (Gem, 2024)) that can generate a large amount of code with little human intervention, faster with high efficiency (Dakhel et al., 2023; Tian et al., 2023).
Despite their usefulness and high potential (Wang et al., 2024; Liang et al., 2024), previous studies have shown that there are many quality (Yetiştiren et al., 2023; Liu et al., 2024b), correctness (Tambon et al., 2025; Vatai et al., 2024; Liu et al., 2024c), security (Pearce et al., 2022; Majdinasab et al., 2024; Fu et al., 2025) and ethical  (Stalnaker et al., 2024) challenges associated with AI-generated code (i.e., code that is primarily generated with AI coding assistance tools). These challenges stress the need to distinguish AI-generated code from other human-written code (Xu et al., 2025; Idialu et al., 2024) for code review and quality assurance purposes (Xu et al., 2025). To this end, we seek to study whether developers self-declare AI-generated code, explore the characteristics of self-declaration comments, and explore how and why developers self-declare (or do not) AI-generated code. We derive our definition of self-declared AI-generated code from the well-established definition of self-admitted technical debt (Potdar and Shihab, 2014): a self-declared AI-generated code is code that has been explicitly acknowledged in the code comments by the developer as automatically generated by AI tools.
A motivating example: We observed in many GitHub repositories that developers often include code comments indicating that certain lines of code or specific files were auto-generated using AI code generation tools. The study of Fu et al. collected a dataset of such code comments to investigate security weaknesses in AI-generated code (Fu et al., 2025). Other studies have also used a similar approach of collecting AI-generated code from GitHub repositories (Yu et al., 2024; Tambon et al., 2025). The authors used code comments from the GitHub repositories to identify AI-generated code. We refer to such comments as self-declaration comments. A sample self-declaration comment from an AI-generated code snippet is illustrated in Figure 1. In this example, a developer acknowledges that the code file was written with the help of GitHub Copilot, an AI code generation tool. This observation motivated us to design a study to explore the practices of self-declaring AI-generated code among developers.
Figure 1. An example of self-declaration comments of AI-generated code
This study aims to empirically investigate the practices of self-declaring AI-generated code. Specifically, we explore the ways developers self-declare AI-generated code and the reasons behind self-declaring (or not) the use of AI-generated code. We address the following Research Questions (RQs):
  • RQ1: How do developers self-declare AI-generated code in their projects?
  • RQ2: Do developers self-declare AI-generated code in practice? Why or why not?
To address our RQs, we utilized a mixed-methods approach combining a repository mining study and a follow-up survey with developers. In the first phase of our study, we identified AI-generated code snippets from open-source projects hosted on GitHub. We collected 613 code files with self-declared AI-generated code snippets from 586 GitHub repositories. We then analyzed and classified those code snippets. Based on the results of the mining study, we conducted a follow-up survey with practitioners to explore the ways developers self-declare AI-generated code and understand the reasons behind their choice to self-declare (or not) their AI-generated code. The survey questionnaire consists of 12 questions, including demographic questions and questions related to self-declaration practices of AI-generated code. Through the survey, we collected valid responses from 111 developers with reasonable diversity (based on location, education, experience, application domains, professional roles, and AI tool preference).
Our findings show that: (1) most developers sometimes self-declare AI-generated code, depending on the proportion of AI-generated code integrated into the project. (2) Developers use various ways to self-declare AI-generated code, ranging from simple self-declaration comments to detailed comments that include the prompt used to generate the code, an explanation of the functionality of the code, and other code quality indications. (3) Developers tend to self-declare their code if the AI-generated code was not substantially modified by humans. When they just use AI-generated code as a guide and then modify the code extensively, they may optionally self-declare AI-generated code. (4) As their reasons for self-declaring AI-generated code, developers reported the need for tracking and monitoring the code for future review and debugging, and also ethical considerations of using AI code generation tools. (5) Developers’ reasons for not self-declaring AI-generated code include the need for extensive human modifications of AI-generated code and the perception that self-declaration is unnecessary, as they view it as comparable to seeking help from online documentation or developer forums.
The main contributions of this work are: (1) a curated dataset of 613 self-declared AI-generated code snippets from open-source repositories on GitHub and performing a comprehensive analysis with a focus on self-declaration comments to understand the practices of self-declaring AI-generated code, (2) an industrial survey to validate the findings of the mining study and explore the potential reasons behind self-declaring (or not) AI-generated code, and (3) guidelines for practitioners to self-declare AI-generated code based on the results of the mining study and the follow-up industrial survey.
Paper Organization: Section 2 discusses related work to this study. Section 3 presents our research questions and outlines the methodology employed for our data collection (from the mining study and the industrial survey). We present our results in Section 4, followed by the discussion of the findings and implications in Section 5. We then present potential threats to the validity in Section 6 followed by the conclusion in Section 7.
2. Related Work
Several studies have looked into different aspects of AI-generated code, in terms of the tools and model used, and their application in software engineering tasks. This includes studies that investigated the usability of AI code generation tools (Liang et al., 2024; Zhou et al., 2025; Barke et al., 2023), quality and correctness (Liu et al., 2024a; Ouyang et al., 2024), and security (Fu et al., 2025; Pearce et al., 2022; Majdinasab et al., 2024). Previous research has also explored the ethical aspects of using AI-generated code in software development (Stalnaker et al., 2024; Xu et al., 2025).
AI Code Generation in Practice. One of the most studied aspects of AI code generation is its use in practice, with several surveys and user-based studies looking into adapting code generation in practice. Zhou et al. investigated GitHub Copilot common problems and solutions from users’ point of view, using data from GitHub issues and discussions and Stack Overflow posts (Zhou et al., 2025). The study identified the issues from users, including compatibility issues, copyright and policy issues, and operational issues. Dakhel et al. studied Copilot by generating solutions to fundamental programming problems and compared them with human solutions (Dakhel et al., 2023). They found that GitHub Copilot can produce solutions to fundamental programming problems, however, those solutions may contain bugs. Barke et al. conducted an observation study with 20 participants while using Copilot in practice. They found that developers use these tools in two different modes: acceleration mode, when they know how to solve the problem but use these tools to solve it quickly, and exploration mode, where they explore various ideas to solve a problem (Barke et al., 2023). Similarly, Liang et al. conducted a large-scale survey of 410 participants on the usability of AI code generation tools (Liang et al., 2024). They found that developers use these tools to enhance productivity by helping them recall syntax and reduce keystrokes. Abrahamsson et al. studied the full-stack development capabilities of ChatGPT instead of generating small code snippets (Abrahamsson et al., 2024). Their findings demonstrate ChatGPT’s ability to create base components, decompose coding tasks, and improve development workflow, with notable improvements in ChatGPT-4 over ChatGPT-3.5, such as improved working memory and more coherent dialogues.
Quality, Correctness and Security of AI-Generated Code. Several recent studies have also investigated the quality and correctness of AI-generated code. Liu et al. evaluated ChatGPT-generated solutions to programming problems and studied their correctness and quality (through static analysis) (Liu et al., 2024a). The results showed that ChatGPT could generate correct solutions almost 70% of the time but suffers from maintainability issues. Ouyang et al. studied ChatGPT code generation capabilities and found that the non-determinism of ChatGPT affects the correctness and consistency of the generated code (Ouyang et al., 2024). Clark et al. evaluated ChatGPT-generated code quality and found that ChatGPT can generate good quality code through iterative prompts, but the generated code was found to contain bugs (Tambon et al., 2025), citing that ChatGPT-generated code would usually require review before integration into projects. These results show that AI-generated code will not hinder the software development process. Bui et al. introduced a framework (called OPENIA) that uses the internal states of code-generation LLMs to evaluate the correctness of AI-generated code (Bui et al., 2025). This approach was found to outperform existing methods with significant improvements in performance metrics, enabling more effective quality assurance in AI code generation.
Similarly, the security of the generated code is another aspect that has been widely studied. Pearce et al. (Pearce et al., 2022) prompted Copilot to generate code in scenarios related to high-risk security weaknesses (CWEs), covering multiple programming languages, and found that 40% of the generated code contained security vulnerabilities. Majdinasab et al. replicated the study by Pearce et al. using Python examples, and found that the newer versions of Copilot improved in terms of the number of security weaknesses in the generated code. However, they still provided vulnerable code suggestions (Majdinasab et al., 2024). Gupta et al. studied ChatGPT from the perspective of cybersecurity, and they found that AI tools such as ChatGPT can be used by both the defenders (secure code generation, vulnerability detection) and attackers (attack payload generation, malware code generation) (Gupta et al., 2023). Fu et al. analyzed AI-generated code from GitHub projects to identify security weaknesses in these code snippets (Fu et al., 2025). The study shows a high likelihood of security weaknesses in AI-generated code found in public projects.
Ethical Concerns with AI-Generated Code. Existing studies have explored ethical concerns with AI-generated code. A recent survey by Stalnaker et al. with developers highlighted the ethical concerns in AI-generated code primarily regarding intellectual property concerns related to copyright violation (Stalnaker et al., 2024). This study shows that developers have mixed opinions about AI-generated code’s copyright and ownership attribution. Xu et al. investigated the ethical concerns and code quality issues with AI-generated code (Xu et al., 2025). These concerns include students’ unethical use of AI-generated code, code plagiarism, and questions of code provenance. The results emphasized the importance of distinguishing AI-generated code from human-written code for code review and quality assurance processes.
Comparative Summary. Existing studies did not study whether developers acknowledge AI-generated code, how they acknowledge it, and what motivates them to acknowledge the use of AI-generated code. We investigate the practices of self-declaring AI-generated code among developers, examine the ways developers use to self-declare AI-generated code, including how they write the self-declaration comments and how they manage its scope, and explore the reasons behind self-declaring AI-generated code. Our study also used a mixed-methods approach to study these practices. Compared to existing work, our study addresses a unique topic with a more comprehensive research methodology. We detail our data collection and research methodology below.
3. Research Methodology
Our study employs a mixed-methods approach comprising two main phases: (1) an analysis of self-declaration comments from AI-generated code collected from open-sourced projects, and (2) an industrial survey with practitioners regarding the practices of self-declaring AI-generated code. We first collected data about AI-generated code from real-world projects by mining GitHub repositories to answer our research question of how developers self-declare AI-generated code. We followed that with a practitioner survey to validate the results from the mining study while addressing the question of why developers self-declare AI-generated code. This methodological triangulation strengthens the validity of our findings by complementing the mining study results with developers’ perspectives from the industrial survey. Figure 2 shows an overview of our research methodology.
3.1. Research Goal and Questions
Our study aims to empirically investigate the practices of self-declaring AI-generated code followed by developers. We are interested in understanding the different ways developers self-declare AI-generated code in their projects, and exploring the developers’ reasons for self-declaring (or not) the integrated AI-generated code. Particularly, we address the following Research Questions (RQs):
RQ1: How do developers self-declare AI-generated code in their projects?
Motivation. Developers may include self-declaration comments in their projects to indicate AI-generated portions of code. These comments may contain important information about the generated code (for example, the tool and prompt used to generate the code). To answer this RQ, we mine GitHub repositories to locate code comments that may indicate if the code is AI-generated. We refer to those comments as self-declaration comments. RQ1 is further divided into five sub-RQs.
RQ1.1: What information do developers share in their self-declaration comments?
This RQ focuses on studying the content of the self-declaration comments collected from the projects. Developers may share useful information about AI-generated code in the added code comments (e.g., the AI tool/model used, prompt to generate the code, additional code explanation, and code quality indications). The information shared in these comments may provide insights into how developers write self-declaration comments.
RQ1.2: What is the scope of the self-declaration comments?
This RQ focuses on analyzing the scopes of the self-declaration comments based on information such as where they are mentioned (i.e., the position of the comment) in the file. The answer to this RQ helps in understanding the common ways developers use to manage the scope of the self-declaration comments, which is valuable for accurately identifying AI-generated code.
RQ1.3: What is the distribution of the AI-generated code snippets with self-declaration comments across repositories?
This RQ focuses on analyzing the frequency of self-declared AI-generated code. We analyze two key aspects: (1) the distribution of the self-declaration comments within individual source files and (2) the number of files containing self-declaration comments within a single repository. This information helps quantify the presence of self-declared AI-generated code within repositories, offering meaningful information on its adoption by developers in practice.
RQ1.4: Are developers satisfied with the AI-generated code according to self-declaration comments?
This RQ aims to investigate developers’ sentiments by categorizing self-declaration comments that contain sentiment information as either positive or negative. A positive self-declaration reflects developers’ satisfaction with the integrated AI-generated code, whereas a negative self-declaration indicates developers’ dissatisfaction with the integrated AI-generated code.
RQ2: Do developers self-declare AI-generated code in practice? Why or why not?
Motivation. AI code generation tools can speed up development and boost developers’ productivity (Ebert and Louridas, 2023; Liang et al., 2024). Developers increasingly use AI code generation tools in practice (SO-, [n. d.]). Here, we aim to investigate whether developers self-declare any AI-generated code used in their projects. We also explore the underlying reasons for developers to self-declare and to not self-declare AI-generated code. The answer to this RQ helps understand the current situation of self-declaration of AI-generated code as a practice in software development. RQ2 is further divided into the following two sub-RQs:
RQ2.1: How common do developers self-declare AI-generated code in their projects?
This RQ investigates how widely developers self-declare (or not) AI-generated code in their projects.
RQ2.2: What are the reasons developers self-declare (or not) the AI-generated code in their projects?
This RQ explores the underlying reasons why developers self-declare (or not) AI-generated code in their projects.
Figure 2. Overview of the mixed-methods research process
3.2. Mining Study Design
To answer RQ1, we collected self-declared AI-generated code from GitHub repositories. We defined a self-declared AI-generated code snippet as one in which the developer explicitly acknowledges that the code was written with the assistance of an AI code generation tool. This self-declaration occurs in the form of code comments added anywhere in the code file. We collected self-declared AI-generated code from GitHub because it is the most widely used platform by developers to host code (Oct, [n. d.]). This allowed us to access code repositories and study projects written in various languages.
We used a keyword-based search to locate code snippets generated by four well-known AI code generation tools: Copilot, ChatGPT, CodeWhisperer, and Code Llama. We prioritized Copilot and ChatGPT due to their popularity and widespread use by developers in practice (Liang et al., 2024; SO-, [n. d.]).
3.2.1. Data Collection
The first step of the data collection process is to identify the search terms to retrieve AI-generated code from GitHub projects. For this purpose, we performed a pilot data search to decide on the most suitable search terms. We used the names of popular AI code generation tools to locate AI-generated code snippets. However, using only the tool names mainly returned the projects where developers shared their experience using AI code generation tools, and did not yield code snippets generated by these tools. We also examined the use of more generic search terms without specifying the name of the tools used (e.g., “AI-generated code”, “AI-assisted code”), which returned only a few AI-generated code snippets. We finally noted that using specific search terms that include the name of the tools (e.g., “generated by GitHub Copilot”, “written by ChatGPT”) can return more relevant results than using those generic search terms. The search terms appreciating AI code generation tools, such as “thanks Copilot” and “thanks ChatGPT” also returned relevant results.
We performed the pilot search using GitHub’s REST API (cod, 2022). As a result, we identified AI-generated code at the repository and code label. The repository and code labels can be referred to as search filters provided by GitHub for easy code searching. We classified the results as being Repository label when the repositories have an explicit acknowledgment of using AI-generated code in the project description (e.g., in the README files). On the other hand, Code label results returned the code files and snippets containing the self-declaration comments about the AI-generated code. The first author classified each code snippet (e.g., whether it is AI-generated), and then the results were cross-validated with two other co-authors. Through this process, we noticed that the search results from the repository label were more likely to contain dummy and student practice projects (i.e., small experimental projects with little content). Considering this, we decided to limit our data collection to the search results from the code label to ensure the quality of our dataset.
In the search results from the code label, we noticed that developers included self-declaration comments in different positions of the code files, acknowledging that the code is AI-generated. The first author collected the code snippets with self-declaration comments from search results that mention the self-declaration of AI-generated code.
# Search Term Results 1 written by ChatGPT 132 2 generated by ChatGPT 111 3 written by Copilot 86 4 thanks ChatGPT 66 5 generated by GitHub Copilot 66 6 thanks Copilot 41 7 by GitHub Copilot 21 8 AI-generated code 19 9 solution by ChatGPT 13 10 from ChatGPT 11 11 help of GitHub Copilot 10 12 solved by GitHub Copilot 6 13 coded by GitHub Copilot 6 14 solution by GitHub Copilot 4 15 from GitHub Copilot 4 16 generated by CodeWhisperer 4 17 from AWS CodeWhisperer 1 18 created by GitHub Copilot 1 19 using GitHub Copilot 1
Table 1. Search terms used for identifying AI-generated code in GitHub repositories
Based on the results of the pilot search, we selected a list of 23 search terms. Of those search terms, 22 included the names of the AI code generation tools and phrases specific to code generation. Only one search term was generic and did not include the name of any specific AI code generation tool. The list of search terms is provided in Table 1, which also shows the number of AI-generated code snippets included in our dataset.
3.2.2. Search and Filtration Process
We applied the finalized search terms to locate relevant code snippets from GitHub projects. The search results were filtered based on the programming languages used in the projects. To increase the language diversity in our dataset, we collected code files from the top 10 programming languages on GitHub  (Top, 2024): Python, Java, JavaScript, TypeScript, C, C++, Go, PHP, Ruby, and C#. Our search process consists of the following steps. First, we used the search terms we identified in Table 1 and applied GitHub’s programming language filter for the top 10 languages. In most cases (19 search terms), the search returned valid results. There were four search terms for which we were unable to retrieve any results that met our inclusion criteria (see Table 2). Specifically, we could not retrieve any results when we searched using the terms “by CodeWhisperer”, “proposed by GitHub Copilot”, “by Llama”, and “from Code Llama”. During the data labeling process, the first author performed the labeling, and then the other co-authors cross-validated the labeling results. We calculated the level of agreement between the authors using Cohen’s Kappa coefficient (Cohen, 1960). The value of Cohen’s Kappa coefficient was 0.82, indicating a high level of agreement among the authors.
Figure 3. Example of included AI-generated code demonstrating inclusion criterion I1
During this process, we applied our inclusion and exclusion criteria (see Table 2). First, we only included code snippets in our dataset when it is clearly stated in the self-declaration comments that they are AI-generated. Secondly, we only included AI-generated code snippets that belong to a repository with more than 1,000 lines of code to avoid dummy and small student practice projects (we used cloc [^1] https://github.com/AlDanial/cloc.git to calculate the lines of code in each repository).
The inclusion or exclusion criteria are provided in Table 2. Figure 3 shows an example of the inclusion criterion I1. This AI-generated code example meets our inclusion criterion 1 because the developer declared in the self-declaration comment that the GenerateMaze() function is generated by GitHub Copilot. It also meets our inclusion criterion 2 because the code is part of the repository GameAwareToys that contains 98 files with 46,198 lines of code, exceeding the 1,000 lines of code threshold. Figure 4 shows an example of the exclusion criterion E1 (see Table 2), as the self-declaration comment did not reveal if the code is AI-generated, but only contained the word “copilot”. Additionally, it belongs to the repository elysia-lambda, which has only 456 lines of code, which is less than the 1000 lines of code threshold (exclusion criterion 2).
Figure 4. Example of excluded AI-generated code demonstrating exclusion criterion E1
Next, we manually checked the source code files for self-declaration comments to determine whether they contain AI-generated code. Once the AI-generated code is identified, we manually reviewed the repositories associated with these code files. To collect a high-quality dataset, we gathered demographic information about the repositories where the AI-generated code was retrieved, including the number of commits, number of contributors, and application domain. We also collected the links to the source code files and self-declaration comments. All of this information was recorded in an MS Excel file, which is included in our dataset (Kashif et al., 2025).
Inclusion Criteria
I1 The self-declaration comment clearly mentions the code as AI-generated.
I2 The source code belongs to a project with more than 1,000 lines of code.
Exclusion Criteria
E1 The self-declaration comment does not clearly indicate that the code is AI-generated.
E2 The code belongs to a project likely to be a dummy or student project with less than 1,000 lines of code.
Table 2. Inclusion and exclusion criteria for mining AI-generated code from GitHub repositories
We categorized the search results into two main categories: snippet-level and file-level self-declarations. Snippet-level is when the self-declaration comment is placed before the code snippet, while file-level is when the self-declaration comment is placed at the beginning or end of the source code file. We also labeled self-declaration comments in which developers’ sentiment regarding the AI-generated code is evident in the self-declaration comments. We manually categorized the comments in which developers appreciated the assistance of AI code generation tools as positive and the comments in which developers shared warnings and disclaimers about the use of AI-generated code as negative. We included all files we analyzed in our dataset (Kashif et al., 2025).
3.2.3. Data Extraction
In the data extraction, we first checked each repository description (README file) to understand the nature of the projects and the application domains. We ensured that we documented the exact line in the code file where the self-declaration comment was mentioned. We extracted the repository name, the line numbers, the number of self-declaration comments instances in a single file, and the total number of files containing AI-generated code snippets in a repository. We also extracted other demographic information, including the number of contributors, the number of commits, and the AI code generation tools used. This recorded information was constantly verified and reviewed by all co-authors. Table 3 provides an overview of our data extraction.
# Data Item Description Related RQ
D1 Self-declaration comments Self-declaration comments that indicate the code as AI generated. RQ1.1, RQ1.4
D2 Repository name The name of the repository to which the code snippet belongs. General
D3 File name File name and direct link to the source code file. General
D4 Application domain Description and the project. Demographic
D5 Programming Language Main programming language of code snippet. Demographic
D6 Code commits and contributors Statistics obtained from the repositories. Demographic
D7 Keyword used Keywords used to retrieve the code snippets. General
D8 Total files with SDC The total files containing self-declaration comments. RQ1.3
D9 Multiple SDC in a single file Extracted multiple self-declaration comments in a single source code file. RQ1.3
D10 Comment type Classification of the comment (snippet-level or file-level). RQ1.2
Table 3. Data items extracted from GitHub repositories
3.2.4. Data Analysis
We employed grounded theory to qualitatively analyze our data (Stol et al., 2016). Grounded theory is a bottom-up approach emphasizing theory generation instead of extending or verifying existing theories (Stol et al., 2016). Specifically, we employed open coding and constant comparison (Stol et al., 2016) to analyze self-declaration comments data collected from GitHub repositories.
Open Coding. We used the open coding technique by identifying and labeling essential ideas and concepts from self-declaration comments. The first author carefully read each self-declaration comment to extract codes. We iteratively refined the code as we examined more data. The other co-authors continuously reviewed the codes and shared their feedback to improve the reliability of the coding process. The analysis was done using MAXQDA [^2] https://www.maxqda.com/ to organize and manage the coding results.
As an example of open coding, we labeled the self-declaration comment, “NB: Much of this was written with Copilot! (although still had to fix a bunch of issues)” as a negative code quality indication. Similarly, we labeled the following self-declaration comment, “create a public function that handles request to the Bitcoin API, this code is generated by GitHub Copilot” as a short code explanation.
Constant Comparison. We developed high-level concepts and categories using the constant comparison method. This was performed iteratively as we compared the codes in individual self-declaration comments (comments under study) with the codes from all self-declaration comments we have encoded so far. The purpose of comparing codes against each other is to group codes with similar meanings into high-level categories. We developed a total of 7 high-level categories from the qualitative analysis of self-declaration comments. For instance, we categorized all positive and negative code quality indications mentioned in self-declaration comments into the code quality indications category. Similarly, we categorized the self-declaration comments containing various code descriptions and explanations into the code explanations category. The first author did the initial coding, while the other two co-authors provided feedback on these concepts by reviewing the coding. We resolved disagreements using the negotiated agreement approach (Campbell et al., 2013) to enhance the reliability of our data analysis.
3.3. Survey Design
The second phase of our study is the practitioner survey, where we investigate how and why developers self-declare their AI-generated code. The survey is motivated by the findings of the mining study, described above in Section 3.2. The survey workflow is shown in Figure 5. We conducted the survey online to allow easy distribution to a vast number of participants. Our questionnaire includes both open-ended and closed-ended questions.
Figure 5. Overview of the survey questionnaire
3.3.1. Questionnaire Design
We designed the questionnaire with a clear focus on our research objectives to ensure that each Survey Question (SQ) is directly aligned with our Research Questions (RQs). The survey begins with an information page that explains the study’s context and outlines the requirements for participating in this survey. The questionnaire is structured into five sections, as follows:
  1. (1)
    Demographic questions: General information about the participants, including educational background, current position, location, and experience.
  2. (2)
    Experience with AI code generation tools: Reflection on developers’ experience and preferences regarding the use of AI code generation tools in practice.
  3. (3)
    Ways of self-declaring AI-generated code: Exploration of the various ways developers use to self-declare their AI-generated code.
  4. (4)
    Reasons for self-declaring AI-generated code: Examination of the motivations behind developers’ decisions to self-declare their AI-generated code.
  5. (5)
    Reasons for not self-declaring AI-generated code: Examination of the motivations behind developers’ decisions not to self-declare AI-generated code.
We provide an overview of our survey questionnaire in Table 4, which is also accessible online [^3] https://forms.gle/WUaFEKBPpfndLfXf8. We distributed the survey through our connections and various mailing lists from early September 2024. The survey was open for participants until mid-October 2024.
ID Section Question Data Analysis Method
Q1 Introduction Do you use AI-generated code in your projects? Descriptive Statistics
Q2 Demographic Which country are you working in? Descriptive Statistics
Q2 Demographic What is your highest educational degree? Descriptive Statistics
Q4 Demographic How many years of experience do you have? Descriptive Statistics
Q5 Demographic What is your role in your organization? Descriptive Statistics
Q6 Demographic
What are the domains or areas of your software projects
in which AI-generated code is declared?
Descriptive Statistics
Q7 RQ1, RQ2 Which AI-code generation tools you use in your projects? Descriptive Statistics
Q8 RQ2
When you use an AI-code generation tool to generate code,
do you self-declare it in your projects?
Descriptive Statistics
Q9 RQ1
If you self-declare AI-generated code, how do you self-declare
them in your projects?
Descriptive Statistics
Q10 RQ1
Why do you use the practice(s) you mentioned in previous
question to self-declare AI-generated code?
Open Coding and Constant Comparison
Q11 RQ2
What are the reasons for self-declaring AI-generated code
in your project?
Open Coding and Constant Comparison
Q12 RQ2 What are the reasons for not self-declaring AI-generated code? Open Coding and Constant Comparison
Table 4. An overview of survey questions and their analysis method
We followed a thorough review and validation process to ensure the survey questionnaire captured the necessary information. We carefully reviewed the questionnaire’s content and structure, and cross-validated its contents to enhance clarity and relevance. We then sought the expertise of an external researcher (non-author) to validate the clarity and rationale of the survey. To that end, we invited a senior researcher with strong expertise in qualitative software engineering research to review the questionnaire for enhancing its clarity. Following his comments, we revised the structure and questions to improve the survey clarity. For instance, we added a prerequisite question to ask participants if they actively use AI code generation tools in practice. We also decomposed the survey into smaller sections to improve readability.
Next, we conducted a pilot run of the survey by distributing the questionnaire to a small subset of our target population. These participants were primarily practitioners with whom we had direct contact in the industry, allowing us to access them for feedback and suggestions easily. We received 11 valid responses from the participants of the pilot survey and made several revisions to the survey questionnaire to improve its structure and readability.
3.3.2. Data Collection
We identified our target population as software engineers who potentially use AI code generation tools in their practice. We distributed the survey to the target population, seeking responses from diverse groups with different experience levels and roles in their organizations. We did not impose any restrictions on participants based on their job titles, geographical location, educational qualifications, or levels of experience.
One challenge we encountered was the difficulty in employing a probabilistic sampling method. It is not feasible to contact every individual within our target population for participation in the survey. As a result, we opted for non-probabilistic sampling methods, which include convenience and snowball sampling. We implemented four main strategies to effectively reach our target population and ensure that our sample accurately reflects the overall population:
  1. (1)
    Recruit practitioners from our contacts who are willing to contribute to our survey utilizing the convenience sampling method.
  2. (2)
    Invite contributors of the repositories we collected during our mining study to participate in our survey using purposive sampling.
  3. (3)
    Request participants to invite practitioners to participate in our survey using a snowball sampling approach.
  4. (4)
    Invite participants through social media platforms such as Facebook and LinkedIn using convenience sampling.
These three strategies were designed to facilitate our data collection process while acknowledging the limitations inherent in non-probabilistic approaches. We sent invitation emails to the contributors of the repositories we collected during our mining study. We invited practitioners in our network via email and LinkedIn messages. We also invited participants through survey invitations posted on our social media profiles and various developer communities and groups on social media platforms, such as LinkedIn and Facebook.
3.3.3. Data Analysis
Validating Responses. To maintain the reliability of our data, we developed a set of inclusion and exclusion criteria to validate each survey response (as shown in Table 5).
Inclusion Criteria
I1 The participant uses AI-generated code in their software projects.
I2 The participant has shared meaningful and consistent responses.
Exclusion Criteria
E1 The participant does not use AI-generated code in their software projects.
E2 The participant has shared meaningless and inconsistent responses.
Table 5. Inclusion and exclusion criteria for collecting survey responses
We included a prerequisite question about whether participants have used AI-generated code in practice. We received 12 responses from practitioners who declared that they do not use any AI-generated code in their projects, and thus, we excluded those responses from our analysis. We also excluded 15 other responses that were unclear or contained irrelevant answers to the survey questions.
Analyzing Responses. We analyzed all responses to the survey questionnaire in Table 4 using a combination of quantitative and qualitative analysis methods. For the first nine closed-ended questions, we employed descriptive statistics to ascertain the frequency of various responses, which provides an overview of the participants’ demographics and their usage of AI-generated code. For the remaining three open-ended questions, we utilized Grounded Theory (Stol et al., 2016), specifically open coding and constant comparative analysis. The first author carefully examined each valid response to extract meaningful insights. During this process, the author summarized the core ideas articulated in participants’ responses regarding their reasons for self-declaration of AI-generated code. For instance, in the answer to SQ11 about reasons for self-declaring AI-generated code, a participant expressed a positive view of self-declaration, stating in $R_{62}$, “It fosters trust with clients and team members by being upfront about the use of AI in the development process”. This response was coded as “fostering trust through transparency”. Another participant emphasized the importance of tracking AI-generated code, stating in $R_{69}$, “Placing comments at the exact location of AI-generated code allows for pinpointed identification of machine-generated contributions, which is essential for precise tracking and review”. This was coded as “tracking AI-generated code for review”. In contrast, some participants provided reasons for not self-declaring the use of AI in development in their answers to SQ12. A participant mentioned in $R_{46}$, “I’ll check the code over myself, the AI is just there to speed things up”, which we coded as “do not rely on AI tools but use them just for assistance”.
In this analysis, we employed constant comparisons to relate and contrast the codes identified from participants’ responses. The first author compared the codes derived from specific summarized ideas to identify patterns and similar semantic meanings. For example, responses that reflected trust, transparency, and accountability were grouped into the transparency, accountability, and ethical considerations category. Responses that emphasized the need for self-declaring AI-generated code to track and monitor it for future code reviews were categorized into the tracking and monitoring code for later reviews and debugging category. All authors were involved in the validation of the developed codes. In cases of disagreement, all authors engaged in discussions to reconcile differences and reach a consensus on the coding results.
4. Results
4.1. Mining Study Overview
We collected 613 AI-generated code snippets from 586 GitHub repositories. We gathered information about the collected code snippets to investigate the nature of the repositories to which they belong, including the programming languages and application domains, and to explore the AI code generation tools used.
Code generation tools. Our study focused on four AI code generation tools: Copilot, ChatGPT, CodeWhisperer, and Code Llama (see details of the selection process in Section 3.2). We found that the highest number of AI-generated code snippets are generated by ChatGPT (336, 56.0%), followed by Copilot (257, 43.0%). The remaining five snippets (¡1%) are generated by CodeWhisperer. We did not find any code snippets generated by Code Llama, as declared by the developers themselves.
Figure 6. Distribution of collected code snippets across programming languages and application domains
Programming languages. A summary of the collected code snippets concerning the programming languages used is shown in Figure 6. We collected self-declared AI-generated code snippets from the 10 most popular programming languages on GitHub. Most of the collected code snippets were written in JavaScript (157, 25.7%), followed by Python (154, 25.2%), Java (68, 11.1%), TypeScript (62, 10.1%, ), C/C++ (59, 9.6%), C# (44, 7.2%), and Go (37, 6.0%).
Application domains. We found self-declared AI-generated code in projects belonging to various application domains, with Web applications being the most common domain with 191 (38.3%) repositories, followed by Utility tools with 162 (32.7%) repositories and Game applications with 79 (15.9%) repositories. 89 repositories were not categorized because the project descriptions were either not provided or unclear. Figure 6 shows a summary of the distribution of self-declared AI-generated code snippets across application domains.
Commit history. Figure 7 shows the commit history of the repositories that contain self-declared AI-generated code. We categorized repositories into six commit count ranges: from 1 – 50 to $\geq$ 5000. In total, we collected data from 586 repositories. 194 (33.1%) repositories are in the 100 to 500 commits range, indicating reasonable activity in these projects. Moreover, 137 repositories (23.4%) contain commit counts ranging from 1 to 50, showing certain inactivity in these projects. The projects containing more than 500 commits account for 136 (23.2%) repositories.
Figure 7. Distribution of repositories based on number of commits
Repository contributors. Of the 586 repositories, we were able to identify the contributors’ information in only 322 of them. The absence of contributor details in the remaining repositories can be attributed to privacy settings (e.g., profiles are not made public) or repositories authored solely by one contributor. As shown in Figure 8, these repositories are maintained mostly by small teams. Specifically, 99 of those repositories have between 2 and 3 contributors, while 88 repositories have between 3 and 5 contributors, indicating that small teams maintain 58.0% of the repositories. Additionally, 64 repositories feature 5 to 10 contributors. Moreover, there are 52 repositories (35.6%) with 10 to 50 contributors. A relatively small proportion of repositories (19, 11.0%) have more than 50 contributors.
Figure 8. Distribution of repositories with multiple contributors
Finding 1 The collected code snippets are primarily written using ChatGPT and GitHub Copilot. Almost half of these code snippets were written in JavaScript and Python, and Web application was the dominant application domain. The repositories that contain the collected code snippets in our dataset are maintained mainly by small teams (less than five contributors, 58.0%).
4.2. Survey Demographics
Our survey covers a diverse sample of participants from industry, with 111 valid responses collected from 26 countries, including participants from Pakistan (40, 36.0%), China (13, 11.7%), Nigeria (8, 7.2%) and New Zealand (6, 5.4%). We received 44 responses from other countries, including Australia, the United States, Bangladesh, Brazil, and the United Kingdom, among others. A summary of participants’ location is shown in Figure 9.
Figure 9. Overview of countries and continents of survey participants
Educational background. We also analyzed the educational background of all survey participants. All participants hold at least a Bachelor’s degree in computer science or related fields. Among the participants, 39 hold a Master’s, while 8 hold doctoral degrees. Figure 10 shows a summary of the participants’ educational backgrounds.
Experience. Participants consist of developers with all levels of experience. Nineteen participants have 1 to 2 years of experience, 28 participants have between 2 and 4 years of experience, 27 participants have 4 to 10 years of experience, and 6 participants have over 20 years of experience. A summary of the level of experience is shown in Figure 10.
Figure 10. Overview of education and experience of survey participants
Professional roles. Most participants (74, 67.2%) reported that they work in software development roles in their organizations, such as software developer, full stack developer, backend developer, and mobile application developer. Participants also work in management positions, including team leads, engineering managers, and CTOs (9 participants), as well as researchers (9 participants). Other roles represented in the survey include quality assurance and testing (5 participants), data-related positions (4 participants, e.g., data analysts and database administrators), security and cloud engineering (3 participants), and design and UI/UX (2 participants). An overview of the participants’ roles in their organizations is shown in Figure 11.
Figure 11. Overview of survey participants’ roles in their organizations
Application domains. The participants work in various application domains. The largest group, comprising 82 participants) works mainly in Web applications, followed by 51 participants who work in Mobile applications, 40 in AI applications, 33 in utility tools, and 12 in games.
AI tools used. The most commonly used AI code generation tool noted by participants was ChatGPT (102), followed by Copilot (52), Claude (28), CodeWhisperer (12), and Gemini (3).
Finding 2 Our survey received 111 valid responses from 26 countries. The survey participants mostly work in the industry in software development-related roles. The majority of survey participants work on web applications and prefer ChatGPT and GitHub Copilot for generating code.
4.3. RQ1: How do developers self-declare AI-generated code in their projects?
4.3.1. RQ1.1: What information do developers share in their self-declaration comments?
We extracted self-declaration comments (i.e., comments associated with the code that helped us identify if the code was AI-generated) from the collected AI-generated code snippets. We present the categories of the self-declaration comments and related examples of such comments in Table 6. These categories show that the common ways developers write self-declaration comments are: 1) simple self-declaration (simple comment mentioning that this code is generated by AI), 2) code explanation (a brief description of AI-generated code), 3) contextual information (context of code generation such as prompts given to the AI code generation tool), and 4) code quality indications (include observations about the quality of the AI-generated code). We discuss each category below. The categories identified in the mining study and the survey study, and their mapping relationships, are presented in Figure 12.
Figure 12. Ways of self-declaration comments from the survey study and their mappings to the ways from the mining study
1) Simple self-declaration. We identified 214 (34.9%) self-declaration comments as simple self-declaration comments. These comments self-declare AI-generated code in a few words, simply acknowledging the code is AI-generated. In most cases (over 96%), the comment will also include the name of the code generation tool used. The example (Comment 1) in Table 6 shows a developer mentioning that Copilot was used to generate the code. Only 3.3% of these comments just self-declared the AI-generated code without mentioning the tool name. For example, we can see in Comment 3 in Table 6 that a developer simply self-declared a function as AI-generated without mentioning any further details. In some self-declaration comments (4%) also acknowledged the simultaneous use of multiple AI code generation tools. In Comment 2 in Table 6, a developer mentioned GitHub Copilot and ChatGPT simultaneously for code generation.
The survey participants also indicated that they used simple self-declaration for AI-generated code. Sixteen participants (14.4%) reported that they simply provided the name of the AI code generation tool used for code generation in the self-declaration comment. Participants mostly shared reasons for using simple self-declaration comments to track and distinguish AI-generated code from human-written code. In $R_{28}$, a participant mentioned that “I only need to distinguish ‘my code’ from AI-generated code”.
Categories Self-Declaration Comment Percentage (%)
Simple Self-declaration 1. This function was mostly written by Copilot (comment) (214) 34.9%
2. Coded by Kristin Witte and GitHub Copilot (and ChatGPT) (comment)
3. Note: AI-generated function (comment)
Code Explanation 4. Function that creates lighter shades of a color, written by Copilot (comment) (185) 30.0%
5. Grayscale the entire page except the specified element solution by ChatGPT (comment)
Contextual Information 6. Function written thanks to ChatGPT (ChatGPT conversation link) (comment) (209) 34.0%
7. Written by Brian Bird, 4/9/2024 with the assistance of GitHub Copilot (comment)
Code Quality Indications 8. Player Class Entirely written by humans with assistance from Copilot (75% human, 25% AI) (comment) (110) 18.0%
9. # NOTE: (11-24-2023) most of this code was generated by ChatGPT (GPT-4). It is a quick solution and needs refinement (comment)
Table 6. Ways and their examples of self-declaration comments from the mining study
2) Code explanation. We also found comments (185, 30.0%) where developers shared the explanation of AI-generated code. These comments explain how the code is generated, what it is about, and how it works. In most cases, these code explanations are brief (83.0% of the 185 comments), typically one sentence long; however, some comments contain more significant details (17.0%). In the example shown in Table 6 (Comment 4), a developer self-declared the function shadeColor(color, percent) as AI-generated code and also explained that its functionality is to generate a lighter shade of a color. Similarly, in Comment 5, a developer self-declared and explained the function grayscale_all_but(element) that is used to grayscale the entire page, except for the element passed as a parameter.
The survey results further confirm these results. We found that 39 participants (35.2%) cited that they provide code explanations when they self-declare their AI-generated code. Participants reported that they typically include text that helps other developers understand the added code. For example, one of the participants explained their reason for writing code explanations in their self-declaration comments in $R_{19}$, “I like to explain the AI-generated code because it helps others understand what it does and why it’s there”.
3) Contextual information. We found 209 (34.0%) self-declaration comments where the details provided additional contextual information about the added code, which includes information such as the author’s name, the date the code was written, the version of the AI tool, and the prompts used to generate the code. For example, Comment 6 shown in Table 6 shows a developer who shared the link to the ChatGPT conversation on generating the function cct2chs(w: number, s: number, v: number). Similarly, we can see that a developer added their username and the date the code was generated in another example (Comment 7 in Table 6).
Several survey participants also expressed a similar view as they also include contextual information about the code when they self-declare AI-generated code. Twenty-one participants (18.9%) mentioned that they provided the prompt to generate the code in the self-declaration comment, with a further 17 participants (15.3%) also reporting that they also provided the date the code was generated.
4) Code quality indications. Several self-declaration comments include a comment about the quality of the generated code, in both positive and negative tones. 48 (8.0%) of comments show that the AI-generated code was modified by a human to transform it for specific requirements. This shows the developer’s contribution to the code, a positive sign for code quality. Some of these comments also mentioned that the generated code was tested and worked per their use case. For instance, in Comment 8 in Table 6, a developer self-declared “player” class and specified that only 25.0% of the code is generated using AI and 75.0% is the human contribution, which is a positive indication. On the contrary, 62 (10.0%) comments highlighted the possibility of errors in the AI-generated code and asked for additional testing and improvements. This type of self-declaration comment shares negative indications about the quality of AI-generated code. Comment 9 in Table 6 shows an example of such comments. In this example, a developer admitted that they provided a quick solution that requires refinements in the self-declaration comment, which is a negative sign for code quality.
Similarly, in our survey, 20 participants (18.0%) reported including a note about the code quality when self-declaring AI-generated code. A participant explained that the reason for including code quality indications in their self-declaration comments is “to keep track of whether the code is modified by a human person and tested according to the requirements of project” ($R_{10}$).
Finding 3 Developers employ various methods to self-declare AI-generated code in their projects, including simply stating the use of AI in a single sentence, providing a brief explanation of the code, sharing contextual information (such as the prompts used to generate the code), or reporting the code quality either positively or negatively.
4.3.2. RQ1.2: What is the scope of the self-declaration comments?
We identified two main self-declared AI-generated code scopes: snippet-level and file-level.
Figure 13. An example of file-level self-declaration comment declaring the entire file as mostly AI-generated
File-level comments. In the file-level self-declaration comments, a developer self-declares the entire file as AI-generated without specifying the exact location within the file. These comments have two possible cases: 1) the entire file is mostly AI-generated with little or no human contribution, or 2) the specific code file contains some AI-generated code, not mentioning the exact location of the code snippet. These comments are typically added at the start or end of the source code file. An example of a file-level comment of the first case is shown in Figure 13. The developer mentioned at the beginning of the file that the added code is mostly AI-generated. Therefore, the scope of this self-declaration comment is the entire source code file. In total, we found 187 (30.5%) file-level self-declaration comments.
The responses to the industrial survey show similar results, with 33 participants (29.7%) indicating that they use file-level comments to self-declare their AI-generated code. A participant explained when they employ file-level comments: “I use file-level comments when I use intensively integrated AI-generated code, and I use snippet-level comments when integrate less AI-generated code” ($R_{10}$).
Snippet-level comments. Developers include these comments at specific locations of the source file (such as before a method/function or a class) to indicate that this particular portion of the code is AI-generated. For example, if a developer adds a comment above a function, it suggests that the comment pertains to that specific function. Figure 14 shows an example of snippet-level comment. In this example, the developer self-declared the AI-generated code before the fixIndentation() function. This suggests that the self-declaration comment applies only to the function. Developers self-declare various code sections in snippet-level comments, including methods/functions, classes, conditions, loops, or statements. Our findings indicate that developers primarily self-declare AI-generated code at the class and function level. We found 426 instances of snippet-level self-declaration comments, accounting for 69.4% of all collected comments.
In the industrial survey, 34 participants (30.6%) mentioned that they utilize snippet-level self-declaration comments to self-declare their AI-generated code, and reported the reasons such as showing exactly which part of the code is AI-generated as shared by $R_{19}$, “I use snippet-level comments to point out the AI-generated parts because it gives clear context right where it’s needed. This way, anyone looking at the code can easily see what was generated by AI”.
Figure 14. An example of snippet-level self-declaration comment declaring a function as AI-generated
Finding 4 Developers use snippet-level comments more often than file-level comments to self-declare AI-generated code. When the code file has a few AI-generated code snippets, developers prefer a snippet-level comment above each AI-generated code snippet. However, if the code is extensively AI-generated, they tend to write a single file-level comment at the beginning or end of the code file.
4.3.3. RQ1.3: What is the distribution of the AI-generated code snippets with self-declaration comments across repositories?
We presented the distributions of AI-generated code in repositories and individual source code files in Figure 15.
Figure 15. Distribution of AI-generated code in repositories and single files
Files containing AI-generated code in a repository. Out of a total of 586 repositories from our dataset, 405 repositories (69.5%) contain only a single file with AI-generated code snippets. We found 178 repositories (30.5%) with multiple files containing AI-generated code snippets, with 139 repositories (23.8%) containing between 2 and 5 files with AI-generated code, while 39 repositories (6.7% ) with more than five files contained AI-generated code.
AI-generated code snippets in a single file. We identified 613 files containing AI-generated code snippets. As presented in Section 4.3.2, most self-declaration comments (426, 69.5%) are at the snippet-level. 400 out of 426 (93.9%) of these files have a single snippet-level self-declaration comment. Out of the 426 files containing snippet-level comments, only 26 (6.1%) have multiple self-declaration comments in a single file. The highest number of self-declarations in a single file is 8 self-declaration comments.
Finding 5 Most of the GitHub repositories in our collected dataset have a few code files with AI-generated code snippets. These code files with AI-generated code snippets typically contain a single self-declaration comment, with only a few code files having multiple self-declaration comments. Overall, the occurrences of self-declaration comments in GitHub repositories are low.
4.3.4. RQ1.4: Are developers satisfied with the AI-generated code according to self-declaration comments?
We analyzed developers’ sentiments concerning their added AI-generated code. This analysis provides a general overview of developers’ satisfaction with using AI-generated code. Through our manual labeling (see Section 3.2), we identified 165 (26.9%) comments that have been labeled as either positive or negative. Overall, we found more positive comments (85.0%) than negative ones (15.0%).
Developer’s satisfaction. In the positive comments, developers acknowledged the usefulness of the AI code generation tools. We provide two examples of such comments with positive sentiments below. In the first example, a developer self-declared a line of code whiteCell = (x // 5) % 2 != (y // 5) % 2 as AI-generated and acknowledged the help of the code generation tool: “Thanks ChatGPT, I was stuck on this for 2 hours” (comment). In the second example, ChatGPT helped the developer to write a CSS code file, which the developer acknowledged positively:“github copilot makes it into a real CSS file for me thanks” (comment).
Developer’s dissatisfaction. We have also found negative comments where developers provided warnings and disclaimers that the code is generated using AI code generation tools. This can indicate that the generated code may contain bugs or may not meet the quality expectations of developers. Some examples of such comments with negative sentiments are: “error handling, really poorly written code, GitHub copilot did it” (comment) and “disclaimer! This code was written by ChatGPT (with minimal editing) and may not necessarily be the best solution for the problem” (comment). Although the developers have integrated these AI-generated code snippets in both examples, they were not quite satisfied with the code quality.
Finding 6 Developers’ sentiments about the AI-generated code and tools used can be identified in some of the added self-declaration comments. The comments are mostly positive, acknowledging the usefulness of the AI code generation tools. However, developers also sometimes express negative sentiments, indicating their dissatisfaction with AI-generated code.
4.4. RQ2: Do developers self-declare AI-generated code in practice? Why or why not?
4.4.1. RQ2.1: How common do developers self-declare AI-generated code in their projects?
The vast majority of the 111 survey participants (90.2%) reported actively using AI code generation tools in their development processes. Since our study focuses explicitly on the practices of self-declaring AI-generated code, we excluded responses (12 in total) where developers indicated that they do not actively use AI code generation tools in practice (see Section 3.3.3).
Figure 16. Distribution of participants based on how frequently they self-declare any AI-generated code
As shown in Figure 16, we see that 63.1% of participants indicated that they sometimes self-declare AI-generated code, 13.5% always self-declare, and the remaining 23.4% never self-declare their added AI-generated code. The self-declaration practices are influenced by the extent to which AI-generated code is integrated into the project. Developers mostly tend to self-declare the use of AI-generated code when it is incorporated with little human modifications. Developers often do not feel the need to self-declare the use of AI-generated code, especially when they use AI-generated code as a guide but then significantly modify the code to meet their requirements.
Finding 7 Most participants revealed that they sometimes self-declare AI-generated code based on the proportion of AI-generated code included in the project and the extent of human modifications to the generated code.
4.4.2. RQ2.2: What are the reasons developers self-declare (or not) the AI-generated code in their projects?
We further explored the reasons behind developers’ decisions to self-declare (or not) their AI-generated code. We present the key reasons we identified from the practitioners’ responses.
Category Answers to Q11 (the reasons for self-declaring AI-generated code) Percentage (%)
Tracking and monitoring code for later reviews and debugging Developers use self-declaration comments to track where AI-generated code is used in the project, allowing them to separate AI-generated and human-written code. This can be helpful for future reference, quality checks, debugging, and improvements. (28) 25.2%
Transparency, accountability, and ethical considerations Developers self-declare AI-generated code for ethical reasons and to maintain transparency and accountability in the development process by clearly stating the origin of AI-generated code. (23) 20.7%
Evaluating and improving code quality Developers self-declare AI-generated code explaining the code quality, which shows the reliability of the code and helps in code improvement in the future, thus improving code readability. (15) 13.5%
Knowledge sharing within the team Developers self-declare AI-generated code to provide context for the code integrated into the project, which enhances knowledge sharing about development within the team and facilitates collaboration. (10) 9%
Table 7. Categories of the reasons for self-declaring AI-generated code (Q11)
Reasons for self-declaring AI-Generated code AI-generated code. Our analysis of survey responses reveals several reasons developers self-declare AI-generated code. We discuss those reasons below (a summary is provided in Table 7).
Tracking and monitoring code for later reviews and debugging. This is the most significant cited reason, as noted by 28 (25.2%) participants. Developers use self-declaration comments to track where the added AI-generated code is used in the project, which can be helpful for future reference, quality checks, debugging, or code review. By clearly self-declaring AI-generated code, developers can more easily trace and address any issues that arise in the future due to the AI-generated code. For example, a participant explained in $R_{9}$ that “by marking exactly where the AI-generated code starts and ends, it gives clear context to anyone reviewing or maintaining the code. They can easily differentiate between AI-generated and manually written parts without searching through the entire file”. Similarly, another participant mentioned that they “use AI-generated code to quickly create blocks of code that would be time-consuming to write. By declaring AI use, if issues arise later, it marks it as an area of code that is worth checking first”.
Transparency, accountability, and ethical considerations. 23 (20.7%) participants noted reasons related to transparency, accountability, and ethics. These participants self-declare the origin of the code to maintain transparency in their teams as they believe that the ethical practice of using AI-generated code is to transparently self-declare it in the project. By including self-declaration comments, developers ensure clarity on which portions of code are AI-generated, promoting responsible use of AI and, thus, promoting accountability. For instance, a participant shared in $R_{108}$ self-declare AI-generated code to ensure transparency and maintain ethical standards, helping collaborators understand its origin and any unique aspects. This practice also supports clearer debugging and proper attribution”. Another participant mentioned in $R_{12}$self-declaring AI-generated code in projects promotes transparency and accountability. It helps users understand the source of the code, encourages responsible use, and allows for better collaboration between human developers and AI tools”.
Evaluating and improving code quality. 15 (13.5%) participants mentioned reasons related to documenting code quality for adding self-declaration comments. Added self-declaration comments are used to explain code quality, indicating the reliability of the code and aiding in future code improvement. For instance, it was noted that self-declaration comments “helps users understand the origin of the code, encourages responsible collaboration, and allows for better evaluation of the code’s quality and reliability” ($R_{27}$). Another participant also cited how self-declaration comments can provide an assessment of the code’s reliability: “it enables assessment of the code’s reliability based on testing and modification history” ($R_{39}$).
Knowledge sharing within the team. 10 (9.0%) participants noted that they self-declare AI-generated code to encourage knowledge sharing within the team. When developers incorporate AI-generated code solutions into their projects, they help other project members understand the code by providing context for the AI-generated portion of the code. This promotes collaboration, helps other team members understand the code, and facilitates knowledge sharing among team members. A participant specifically emphasized sharing the prompts used to add context to the AI-generated code: “including the original prompt used to generate the AI code ensures context is provided for why the code was generated in that specific way” ($R_{71}$). Similarly, another participant reported that “the reason for self-declaring AI-generated code is to facilitate collaboration by providing context for other developers” ($R_{40}$).
Finding 8 Most participants noted that they self-declare AI-generated code because of the need to track it for future code review and debugging due to concerns about code quality. Another reason that was mentioned was the promotion of the ethical use of AI-generated code.
Reasons for not self-declaring AI-generated code. Our survey responses analysis reveals several reasons explaining why developers may choose not to self-declare their AI-generated code. We summarize those reasons in Table 8, and describe them in detail below.
Customization and modification. One of the most frequently cited reasons, as noted by 13 (11.7%) participants, for not self-declaring AI-generated code is that developers mostly edit and thoroughly modify the generated code according to their specific use cases. It is believed that the AI-generated code may not fully address the business requirements or may contain inaccuracies. For these reasons, developers extensively verify, modify, and test the code. Since the code has undergone significant changes, developers feel that it has become their own contribution. For instance, a participant noted that he/she thoroughly modified the AI-generated code before committing it: “The AI-generated code that I use in my application is thoroughly understood, tested and, many times, radically changed before being committed, essentially making it original code” ($R_{93}$). Another participant pointed out that the AI-generated code does not fulfill requirements and emphasized review and modification in $R_{33}$-generated code does not fully meet business requirements and needs to be reviewed and modified”.
Category Answers to Q12 (reasons for not self-declaring AI-generated code) Percentage (%)
Customization and modification Developers do not self-declare AI-generated code because it contains flaws and does not meet the business requirements. They thoroughly review, change, and test it before adding it to their projects, which makes it their own contribution. (13) 11.7%
Self-declaration is unnecessary Developers do not self-declare because they use AI code generation tools to enhance their productivity but do not completely rely on them in their development. They see it as similar to seeking help from online documentation/discussion forums (Stack Overflow) and do not feel the need to self-declare it. (13) 11.7%
Other reasons Developers also shared some less popular reasons, such as laziness, lack of proper requirements from the organizations, or sometimes discouragement from organizations due to security and legitimacy concerns with AI-generated code. (6) 5.4%
Table 8. Categories of the reasons for not self-declaring AI-generated code (Q12)
Self-declaration is unnecessary. The other top-cited reason, as mentioned by 13 (11.7%) participants, for not self-declaring AI-generated code is that some developers see it as an unnecessary activity. It was noted that developers may utilize AI code generation tools to generate ideas; however, they emphasized that they only employ them to enhance their productivity. They see AI-generated code as similar to seeking help from online documentation or developer forums (such as Stack Overflow). Those AI code generation tools are seen as assisting tools. Hence, they believe that self-declaration of AI-generated code is an unnecessary activity. For example, a participant answered that: “I usually search for small code snippets in languages I don’t remember the syntax for, and I believe that adding comments specifying the source would be unnecessary for my team” ($R_{95}$). Another participant noted the following: “In the rare cases where I’m not providing my own original code, it’s going through multiple iterations of my prompting and manually editing to get what I need ….. I don’t self-declare using a spell-checker, so I don’t see the value in self-declaring AI-generated code”.
Finding 9 Participants noted that they do not self-declare AI-generated code because the code requires human review and modification before integration into their projects, and they feel that self-declaration is not required, as it is similar to using online documentation or tools like a spell checker.
5. Discussion
5.1. Developers Practices in Declaring AI-Generated Code
Our study findings demonstrate that developers follow different ways to self-declare AI-generated code in their projects. Developers include varying levels of detail in the self-declaration comments, ranging from simple one-liners to detailed explanations and contextual information about the added code. Developers’ approaches to the scope of self-declaration comments also vary - sometimes self-declaring specific code snippets or entire code files as AI-generated. Developers utilize these ways not only to provide simple annotations but also to address ethical, collaborative, and quality-related concerns. Below, we discuss the key reasons why developers use various ways to self-declare AI-generated code.
5.1.1. Information Developers Mention in the Self-Declaration Comments
We classified the types of information provided by developers in the self-declaration comments, as shown below.
Simple self-declaration. Developers state in the self-declaration comments that the specific file or code snippet is AI-generated. Developers typically share the name of the tool used to generate the code, as well as the name of the code snippet generated (e.g., class or method name). This declaration approach is useful for code traceability as it separates AI-generated code from human-written code. Developers prefer writing simple self-declaration comments to maintain traceability and transparency when integrating AI-generated code while also focusing on efficiency when writing self-declaration comments to avoid complex workflows. Yu et al. also found that comments acknowledging the use of AI are mostly short and lack details (Yu et al., 2024). Simple self-declaration comments ensure that the development process remains transparent and adheres to ethical considerations. It also preserves the code provenance by documenting the origin and history of generated code.
Code explanation. Explanations of code are seen as useful, especially when the AI-generated code is relatively complex. Developers mostly write such comments to explain three aspects of the code: (1) its functionality, (2) how it was generated, and (3) how the code was added. By adding these comments, developers aim to enhance the comprehensibility and readability of the code, a goal also discussed in related studies (Pascarella et al., 2019; Nielebock et al., 2019; Scalabrino et al., 2016). Adding code explanations is particularly valuable for complex or critical code segments, where understanding the rationale behind the code is essential for debugging and future modifications. This way of self-declaration was also widely supported by the survey participants. However, the challenge with this self-declaration approach is to maintain a balance in the amount of details to be shared. This is because overly verbose explanations may hinder readability, while insufficient detail may fail to provide a meaningful explanation of the added AI-generated code.
Contextual information. Developers may also add contextual information in the self-declaration comments. For example, we noted several instances of comments that included the prompt used to generate the code, which can help inform other team members about the initial requirement for which the code was generated. This improves the team’s understanding of the code and its context. Developers also share other information about the code, including the date when the code was generated and the developer’s username for better tracking of AI-generated code. These results are also evident in our industrial survey, confirming that developers share contextual information when self-declaring AI-generated code. Using this practice enables developers to enhance knowledge sharing and collaboration within their teams. The study by Stalnaker et al. shows similar results, where developers mentioned that they record the prompts or note the whole interaction with AI tools to document AI-generated code (Stalnaker et al., 2024). Similarly, Yu et al. reported that developers share the prompts used to generate code (Yu et al., 2024).
Code quality indications. Self-declaration comments can contain code quality indications. This demonstrates developers’ awareness of the code quality risks associated with the use of AI-generated code. Developers often claim that they have integrated AI-generated code after human modification and testing before its use in their projects. On the contrary, they may also warn that the code offers a quick fix with less human intervention and inadequate testing. In both cases, developers mention details about the code quality to provide helpful information on the reliability of the integrated AI-generated code. This way of self-declaration is beneficial for addressing quality issues in AI-generated code, for example, AI-generated code may suffer from security issues (Fu et al., 2025) as well as bugs (Tambon et al., 2025).
5.1.2. Scope of Self-Declaration Comments
We categorized the scope of self-declaration comments into two categories: snippet-level and file-level self-declaration comments. We discuss both categories and their comparison below.
File-level comments. When developers add multiple AI-generated code snippets in one source file, or the source file is mainly written with assistance from AI code generation tools, they provide a single self-declaration comment at the start or end of the source file to acknowledge that the file contains AI-generated code. The drawback of file-level comments is that they are provided at a high level and can not help accurately track the AI-generated code (i.e., which part of the file is AI-generated).
Snippet-level comments. When developers use short AI-generated code snippets in their projects, they mention precisely which specific segments are AI-generated. The self-declaration comments only apply to the specific scope of the self-declared AI-generated code snippets. For example, if a developer writes a snippet-level comment above a function, then the scope of this comment applies only to that specific function. The main benefit of snippet-level self-declaration is that it accurately describes which code snippet is AI-generated, allowing for the separation of AI-generated from the human-written code within a code file. This helps in tracking the AI-generated code in later code reviews, bug fixing, and quality improvement processes.
Snippet-level vs. file-level self-declaration comments. The preference between snippet-level and file-level self-declaration comments reflects developers’ trade-offs between granularity and efficiency. Developers prefer snippet-level self-declaration comments when they need to track and differentiate AI-generated code accurately. On the other hand, developers use file-level self-declaration comments when the code file contains multiple AI-generated code snippets or is predominantly AI-generated. File-level comments are efficient as they save time and effort, but sacrifice granularity, making it difficult to trace individual AI-generated code snippets within code files. Developers’ preference for snippet-level comments, as indicated by the industrial survey and mining study, demonstrates that they prioritize accurate traceability and accountability over efficiency when integrating AI-generated code in their projects.
5.1.3. Distribution of AI-Generated Code.
The distribution of self-declared AI-generated code across repositories (shown in Figure 15) suggests two possible interpretations: (1) Even with the introduction of AI code generation tools in software development, developers still use these tools in a limited manner and do not rely on them due to code quality (Nikolaidis et al., 2024; Dakhel et al., 2023) and security concerns (Fu et al., 2025). The low adoption rate also indicates that developers are still exploring the potential of AI tools and adding them selectively into their development workflows. (2) The actual distribution of AI-generated code in these projects may be higher than reported, as our analysis relies solely on self-declaration of AI-generated code by the developers. Some developers may integrate AI-generated code without declaration or acknowledgment. In our survey, some participants mentioned that they add AI-generated code without self-declaration (see the discussion in Section 5.2.2).
5.1.4. Developer Satisfaction.
By analyzing developers’ sentiments expressed in their self-declaration comments, we found that most developers were satisfied and reported a positive experience in generating the desired AI-generated code. These results align with previous findings, which show that AI code generation tools enhance developers’ productivity (Peng et al., 2023) and are useful in coding tasks (Nikolaidis et al., 2024). These tools enable developers to focus on higher-level problem-solving tasks, as they reduce the time and effort required for development. However, this satisfaction is related to the ability to control and refine the outputs of code generation tools, as highlighted by the emphasis on iterative prompting and human intervention.
5.2. Reasons for Self-Declaring (or Not) AI-Generated Code
Developers shared mixed opinions about their reasons for self-declaring their AI-generated code. We discuss those reasons below.
5.2.1. Reasons for Self-Declaring AI-Generated Code.
Our survey results show that developers self-declare AI-generated code due to several reasons, including (1) the need for tracking and monitoring AI-generated code in code files, (2) promoting ethical use of AI-generated code, (3) providing context to share knowledge about AI-generated code and enhance collaboration, and (4) indicating the code quality by sharing whether the code is tested or undergone human modification.
Tracking and monitoring code for later reviews and debugging. A key reason developers self-declare AI-generated code is to track its usage for future code reviews, debugging, and enhancement. The self-declaration is primarily used to differentiate between AI-generated and human-written code. This distinction becomes particularly important when developers do not verify the functionality or enhance the quality of AI-generated code. There is a need for reviewing and modifying any AI-generated code, with several previous studies denoting possible quality issues in AI-generated code and suggesting continued verification and extensive testing (Jin et al., 2024; Ouyang et al., 2024).
Transparency, accountability, and ethical considerations. Another reason for self-declaring AI-generated code is that it is considered an ethical obligation. Several previous studies have explored ethics in AI, specifically in software development (Stalnaker et al., 2024; Xu et al., 2025). Nonetheless, existing research lacks clear guidelines on the ethical use of AI-generated code in practice. Developers noted that self-declaring AI-generated code facilitates transparency in the development process, and helps track its provenance (Xu et al., 2025).
Evaluating and improving code quality. Self-declaring AI-generated code allows developers to indicate whether the code is entirely AI-generated, tested, or contains human contributions. The self-declaration comments show the reliability of the code by allowing developers to assess the quality of the code and work on reviewing and improving it. This way of writing self-declaration comments aligns with previous studies, which raise concerns about the quality of AI-generated code (Liu et al., 2024a; Ouyang et al., 2024). Due to these concerns, human review and improvements are inevitable for AI-generated code to satisfy the code quality requirement (Dakhel et al., 2023).
Knowledge sharing within a team. Self-declaring AI-generated code helps developers improve knowledge sharing and collaboration by providing explanations and contextual information in the self-declaration comments. In our mining study of GitHub, we identified short code explanations and contextual information, such as prompts used to generate code. Developers include such information in their self-declaration comments to provide context for AI-generated code to other team members, sharing knowledge about the code and enhancing collaboration. Similar to our study, Yu et al. reported that developers share prompts used to generate code and sometimes provide additional details in code comments (Yu et al., 2024). Stalnaker et al. also found that developers record the prompts or the entire interaction with AI tools to document AI-generated code (Stalnaker et al., 2024).
5.2.2. Reasons for Not Self-Declaring AI-Generated Code.
Our survey results show that there are two reasons for not self-declaring AI-generated code: (1) the extensive modifications to AI-generated code and (2) the perception of AI code generation tools as productivity enhancers rather than autonomous contributors.
Customization and modification. The main reason is that developers do not self-declare AI-generated code because the code undergoes a cycle of iterative refinement before it is added to their projects. The rationale behind this modification is twofold: (1) Developers use the generated code as a source of inspiration only. Although AI code generation tools have undergone significant improvements over the years, their output often still requires customization to meet project requirements. (2) AI-generated code may contain code quality issues that require human modification before it becomes a viable solution. As developers extensively verify and modify the code according to their requirements, the generated code evolves through additional human contributions. Therefore, developers feel that the code is now their contribution.
Self-declaration is unnecessary. Another reason is that developers do see those AI code generation tools similarly to reference materials such as online documentation and coding forums (e.g., Stack Overflow). Therefore, they view the generated code as similar to adapting examples from online forums or documentation. Developers utilize AI code generation tools to enhance productivity; however, they do not employ these tools as autonomous contributors. Therefore, they do not feel the need to self-declare AI-generated code.
5.3. Implications
Based on our results, we offer a set of guidelines for developers regarding how and when to declare AI-generated code. This includes recommendations on what information developers should include in their self-declaration comments and how to effectively manage the scope of self-declaration comments.
Self-declaring AI-generated code, especially when the generated code is not fully understood or tested. Developers should self-declare AI-generated code, particularly when it has been generated with little human intervention, or when the generated code has not been thoroughly tested or verified. Even though the code is thoroughly reviewed, understood, and tested, self-declaration remains valuable for the ethical use of AI-generated code. In both cases, self-declaring the code is beneficial for maintaining a record of its origin and supporting future maintenance.
Using simple self-declaration comments for better transparency. When adding AI-generated code to a project, developers may include a short self-declaration comment. This comment should note that the code was AI-generated, the name of the tool used, and the name of the code snippet (e.g., the name of the class, function, or language construct). This practice helps ensure transparency in development, promotes the ethical use of AI-generated code, and facilitates the distinction between AI-generated and human-written code.
Adding code explanations for complex AI-generated code. A good practice for developers is to include a brief explanation alongside the simple self-declaration comments, especially when the AI-generated code is relatively complex. These comments should describe how the code works, why it is added, and any key details about the integrated AI-generated code that team members may need in the future. These code explanations should be kept concise, and developers should be mindful of the code’s readability.
Include code quality indications in self-declaration comments. Developers should add details about the quality of AI-generated code in the self-declaration comments. This can include whether the code has been tested, whether it was modified by a human, and if there are any known issues or limitations in the AI-generated code. Including such information helps team members assess the code’s reliability and determine if further improvements are necessary.
Document the context of AI-generated code. When using AI-generated code, developers should consider documenting its use context, such as the prompts used. Keeping a record of such information can help developers understand how the code was generated and make it easier to modify or troubleshoot in the future.
Using snippet-level comments for better traceability. It is advised to use snippet-level rather than file-level comments when self-declaring AI-generated code, as they provide a precise and more detailed approach to track AI-generated code snippets. However, if a file contains mostly AI-generated code or multiple AI-generated code snippets, file-level comments may be more suitable, as they offer a more concise summary without excessive repetition.
6. Threats to Validity
6.1. Internal Validity
Possible bias in manual analysis. During the data collection of the mining study, we first searched for AI-generated code on GitHub repositories using a variety of search terms. After we got the results, we manually labeled the code snippets as AI-generated. This approach may introduce bias in our manual labeling of AI-generated code snippets. To reduce this bias, the first author performed the labeling, and then the other co-authors cross-validated the labeling results. In cases of disagreements, we performed the labeling process through extensive discussion and continually measured the agreement level (Cohen’s Kappa coefficient) among the authors.
Participant familiarity with AI-generated code. Our study focuses specifically on AI-generated code, and we wanted survey participants who actively use AI code generation tools in their projects. This requirement introduces a threat that some respondents may lack experience with AI code generation tools or may prefer not to use them in their software projects. To minimize this risk, we included a prerequisite question on the first page of the survey that asks participants if they use AI-generated code in their software projects. Based on their answers, we excluded 12 responses from participants who did not use AI-generated code in their software projects. We also clarified that we invite participants who utilize AI-generated code in their projects on the first page of our survey questionnaire.
Selection of participants. We identified survey participants from GitHub repository contributors, our personal contacts, and developer groups, and invited them via email and social media platforms. A potential threat to validity arises if some participants lack experience in software development, making them ineligible to participate in our survey. To minimize this threat, we explicitly mentioned on the first page of our survey that we invite participants who are professionally working in the software industry. In addition, we also included questions about education (survey question Q3) and experience (survey question Q4) in the demographic questions, helping us to ensure that we only received responses from participants with the appropriate software development background.
6.2. Construct Validity
Keyword-based search. We used a set of search terms to identify and locate AI-generated code on GitHub repositories. This approach introduced a threat to validity as the search terms used might not be comprehensive or inclusive. To minimize this threat, we first conducted a pilot search using a small set of keywords, which helped us finalize the list of potential search terms. We iteratively added new terms to improve the coverage of our keyword. As a result, we identified a list of 23 search terms with a broad coverage.
Understandability of the survey questionnaire. We created our survey questionnaire with a total of 12 questions, including nine closed-ended and three open-ended questions. The survey questions may introduce a threat to construct validity. To overcome this threat, we employed two strategies: (1) We requested the assistance of an experienced software engineer researcher to review the survey questionnaire, which helped us refine the clarity of the survey questions and the structure of the questionnaire. (2) We conducted a pilot run of the survey with a small subset of our target survey population. We asked the participants in the pilot survey to provide feedback on the clarity of the survey. We used this feedback to improve the readability and understandability of our survey questionnaire.
6.3. External Validity
Generalizability of the survey results. Our survey received 111 valid responses from practitioners. Although the survey provides important information about the practices of self-declaring AI-generated code, the sample of participants can be viewed as not fully representative. To minimize this threat, we aimed to collect responses from diverse countries and regions to improve the generalizability of our findings. In addition, the diversity of the survey participants, including their experience, roles within the current organization, and application domains (see Section 4.2), partially mitigates this threat.
Collecting code snippets only from GitHub. We collected AI-generated code snippets from GitHub repositories. Considering that AI-generated code is also available on other coding platforms, this poses a threat to the external validity of our study results. To minimize this threat, we compiled a dataset of 613 code snippets from repositories across the top 10 programming languages, ensuring a broad coverage. Given GitHub’s position as the leading open-source data source and its common use in similar research, this threat is reasonably mitigated.
6.4. Reliability
The reliability of empirical studies refers to the ability to replicate results and findings under similar conditions. A potential threat to reliability arises from the manual collection and analysis of AI-generated code snippets from GitHub projects and the qualitative analysis of the survey responses. To mitigate this threat, we followed a rigorous process for collecting and analyzing data from both sources (see Section 3.2 and Section 3.3). We also made the dataset used in the study (Kashif et al., 2025) available for other researchers to replicate our findings.
7. Conclusions and Future Work
This study explores the practices of developers’ self-declaration of AI-generated code. We examined whether developers self-declare AI-generated code, their motivations for doing so, and the methods they use to achieve this. We employed a mixed-methods approach, combining data from GitHub repositories with a follow-up industrial survey among practitioners. We found that developers have varying opinions on self-declaring AI-generated code, and their self-declaration practices are influenced by the proportion of AI contribution in the generated code. Most practitioners (76.6%) always or sometimes self-declare AI-generated code. In contrast, other practitioners (23.4%) noted that they never self-declare AI-generated code. Top reasons for self-declaring AI-generated code include (1) the need to track and monitor code for future review and debugging, and (2) ethical considerations related to the use of code generation tools. On the other hand, reasons for not self-declaring AI-generated code include (1) AI-generated code often requires extensive human modification, and (2) developers view self-declaration as an unnecessary activity, comparable to utilizing online documentation or coding forums. We then proposed a set of guidelines for developers to follow when self-declaring AI-generated code.
Our study also opens up several directions for future research. (1) We identify various reasons for self-declaring (or not) AI-generated code, which can be further investigated with a larger set of developers. Additionally, we provide guidelines for self-declaring AI-generated code by analyzing how developers report their generated code based on data from the mining study and industrial survey. Future research can further refine and expand the guidelines using new datasets and validate their usefulness through empirical evaluation. (2) Our research shows that developers consider self-declaration comments beneficial for tracking AI-generated code for future review, debugging, and improvement. This opens up avenues for future research to investigate the impact of self-declaration comments on various code quality attributes, such as readability and maintainability. Such studies can contribute to a deeper understanding of the evolving landscape of software development, where AI-generated code is increasingly integrated into modern software systems. (3) It would be valuable to provide a code-commenting function in AI code generation tools to support the guidelines (see Section 5.3) for developers to declare AI-generated code.
Data Availability
The dataset used in this work is available at (Kashif et al., 2025).
Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.